I am having difficulty understanding how to find the slope and y intercept of this problem, could you please show me how in step form? Thank you
2x + 3y = 1

Answer by rfer(16322)

use the slope intercept formula
put this equation into that form.
y=mx+b
2x+3y=1
3y=-2x+1
y=-2/3x+1/3
m=slope=-2/3
three right, two down on graph(rise over run)
b=y int=1/3
y int is when x=0
(0,1/3)
Does that help?
Bob